OpenStudy (anonymous):

-3q to the second power - 10q = q to the second power-15+5q

myininaya (myininaya):

\[-3q^2-10q=q^2-15+5q\]

myininaya (myininaya):

get everything to one side

OpenStudy (saifoo.khan):

okay!

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

what is the answer?

myininaya (myininaya):

get everything to one side of the equation

myininaya (myininaya):

let me know what you get when you do this

myininaya (myininaya):

i assume you want to solve for q right?

OpenStudy (anonymous):

-3q^5q-10^15^+5q

myininaya (myininaya):

\[-3q^2-q^2-10q-5q+15=0\]

myininaya (myininaya):

\[-4q^2-15q+15=0\]

OpenStudy (saifoo.khan):

C this/

OpenStudy (anonymous):

which one is it??? im sooooooooo confused

myininaya (myininaya):

now we use the quadratic formula \[q=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

no just combining

myininaya (myininaya):

i gotta go

OpenStudy (saifoo.khan):

rose, post the probs.. One by One.

OpenStudy (anonymous):

-3q to the second-10q +qto the second- 15+5q

OpenStudy (anonymous):

now you gave me a different answer then her i dont need steps I just need answer

OpenStudy (anonymous):

ok. thanx i have a couple more

OpenStudy (saifoo.khan):

ok.

OpenStudy (anonymous):

1/3b - 4 +2/9b-4

OpenStudy (saifoo.khan):

post them seperately

OpenStudy (anonymous):

1/3b -4 +2/9b -4

OpenStudy (saifoo.khan):

use brackets please! :'(

OpenStudy (anonymous):

(1/3b)-4(+2/9b)-4

OpenStudy (anonymous):

4/9b-8?

OpenStudy (saifoo.khan):

yes/

OpenStudy (anonymous):

9.7-8.8b-3.2

OpenStudy (saifoo.khan):

-8.8b +6.5

OpenStudy (anonymous):

A sedan travels 6mph slower than a sports car . write an expression for the speed of the sedan if the sports car travels v mph?

OpenStudy (saifoo.khan):

v-6

OpenStudy (anonymous):

evaluate the expression when x=-2 y= 2/3 z=4 w= -1/2

OpenStudy (saifoo.khan):

the ques is incomlepete.